OpenStudy (anonymous):

an angle that measure 2 pi/3 is a reflex angle true or flase

OpenStudy (espex):

A reflex angle is an angle that is greater than 180 degrees and less than 360 degrees. The given data, 2 pi/3 suggests the angle measured in radians. 1deg = radians * 180/pi \[\frac{2\pi}{3} \times \frac{180}{\pi} \rightarrow \frac{2}{3} \times 180 \rightarrow 2\times 60=120 \space degrees\]